OpenStudy (aglna):

What is the volume of a sphere with a radius of 3 inches

MARC (marc_d):

\[volume~of~sphere=\frac{ 4 }{ 3 }\pi~r^2\]where r= radius

OpenStudy (aglna):

My answer is 12pi the answers it could be are 12pi cubic inches 24pi cubic inches 36pi cubic inches 48 pic cubic inches

MARC (marc_d):

plug in the value to find the volume of the sphere..and leave the value in terms of pi..

MARC (marc_d):

yep,that's right! @Aglna

OpenStudy (aglna):

Thanks

MARC (marc_d):

yw!